# Geometric Mechanics of Thin Periodic Surfaces

### v1

@@ III.2 Shape-Periodic deformations
$E$ and $H$ are the changes to the macroscopic (i.e., averaged over the unit cells) first and second fundamental forms, respectively. In SI Sec. III B 4, we further discuss the physical meaning of each term in Eq. (3.4) and explicitly show that the equation satisfies the formal definition of shape-periodic deformations [Eq. (3.3)] to linear order.
@@ IV Origami and triangulation of smooth surfaces
Figure 4: Numerically computed isometries of surfaces. (a) The surfaces and their triangulations used in the simulations: a Miura-ori (first row), a graph of translation (second row) and a perturbed graph of translation (third row). Explicit parametrizations of the latter two surfaces are provided in Fig. 2 of the SI. Unit cell geometries are shown in the corresponding insets. The unit cell of the triangulated Miura-ori comprises eight triangular faces, whereas each of the other two triangulated surfaces has $2\mathcal{N}^{2}=288$ triangular faces in its unit cell, where $\mathcal{N}$ denotes the number of grid squares per row (and per column) and serves as the parameter controlling the triangulation density. (b) The three numerically computed linear isometries are shown in the three columns, along with their vector representations [of the form in Eq. (3.7)] and the corresponding macroscopic plate deformations, displayed in the insets.
@@ VI ACKNOWLEDGMENTS
The authors acknowledge financial support from the Army Research Office through the MURI program (#W911NF2210219) and through the National Science Foundation CAREER program (#2338492), NSF CMMI- 2247095 (CDS) and through NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). We would also like to thank Evan Dickson for fabricating the 3D-printed surface and folding the Miura-ori shown in Fig. 1.

### v2

@@ III.2 Shape-Periodic deformations
Given the form of a shape-periodic mode $\delta_{\mathrm{SP}}{\mathbf{x}}\left(u^{1},u^{2}\right)$, the corresponding changes in the macroscopic (i.e., unit-cell-averaged) first and second fundamental forms, $E_{\alpha\beta}$ and $H_{\alpha\beta}$, can be obtained using:
$H_{\alpha\beta}$ $=\left\langle{{\mathfrak{d}}_{\alpha}{{\mathfrak{d}}_{\beta}{\delta_{\mathrm{SP}}{\mathbf{x}}}},\hat{\mathbf{z}}}\right\rangle$, (3.5, a)
$E_{\alpha\beta}$ $=\frac{1}{2}\left(\left\langle{{\mathfrak{d}}_{\alpha}{{\delta_{\mathrm{SP}}{\mathbf{x}}}},\hat{\mathbf{\bm{\ell}}}_{\beta}}\right\rangle+\left\langle{{{\mathfrak{d}}_{\beta}{\delta_{\mathrm{SP}}{\mathbf{x}}}},\hat{\mathbf{\bm{\ell}}}_{\alpha}}\right\rangle\right)+H_{\alpha\beta}\,x_{z},$ (3.5, b)
where the lattice derivative of a function $f\left(u^{1},u^{2}\right)$ is defined as ${\mathfrak{d}}_{1}{f}\equiv\left[{f}\left(u^{1}+\ell^{1},u^{2}\right)-{f}\left(u^{1},u^{2}\right)\right]/\ell^{1}$ (and similarly for ${\mathfrak{d}}_{2}{f}$). We note that Eqs. (III.2) exactly mirror the usual expressions used to compute the microscopic infinitesimal strains and curvature changes. [See SI Sec. VI B for a discussion of the gauge-invariance properties of the macroscopic strains $E_{\alpha\beta}$, which give rise to the term $H_{\alpha\beta}\,x_{z}$ in Eq. (3.5, b).] In SI Sec. III B 4, we further discuss the physical meaning of each term in Eq. (3.4) and explicitly show that the equation satisfies the formal definition of shape-periodic deformations [Eq. (3.3)] to linear order.
@@ IV Origami and triangulation of smooth surfaces
Figure 4: Numerically computed isometries of surfaces. (a) The surfaces and their triangulations used in the simulations: a Miura-ori (first row), a graph of translation (second row) and a perturbed graph of translation (third row). Explicit parametrizations of the latter two surfaces are provided in Fig. 2 of the SI. Unit cell geometries are shown in the corresponding insets. The unit cell of the triangulated Miura-ori comprises eight triangular faces, whereas each of the other two triangulated surfaces has $2\mathcal{N}^{2}=288$ triangular faces in its unit cell, where $\mathcal{N}$ denotes the number of grid squares per row (and per column) and serves as the parameter controlling the triangulation density. (b) The three numerically computed linear isometries are shown in the three columns, along with their vector representations [of the form in Eq. (3.8)] and the corresponding macroscopic plate deformations, displayed in the insets. The macroscopic strains and curvature changes associated with each mode are computed using Eqs. (III.2), with the surface average heights set to zero, as shown for example in Fig. 3 (a). (See SI Sec. VII B 3 for more details.)
@@
In our simulations, after triangulating the selected surfaces, we obtain their isometry data—the dihedral angle changes and, subsequently, the isometric displacements—by computing the zero modes of their associated compatibility matrices (see SI Sec. VII B 2 for details). With the isometry data in hand, Eqs. (III.2) are used to extract the macroscopic strains and curvature changes corresponding to the isometries. A caveat applies when we report the numbers: Because the macroscopic strains are gauge-dependent (see SI Sec. VI B for details), the average heights of the selected surfaces are set to zero to produce physically intuitive results; in particular, in this height gauge, bending a surface leaves its mid-height plane strainless.
@@ VI ACKNOWLEDGMENTS
The authors acknowledge financial support from the Army Research Office through the MURI program #W911NF2210219 (DZR) and through the National Science Foundation CAREER program #2338492 (DZR), NSF CMMI- 2247095 (CDS) and through NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP) (DZR). We would also like to thank Evan Dickson for fabricating the 3D-printed surface and folding the Miura-ori shown in Fig. 1.